OpenStudy (anonymous):

X^2(√(1+x))dx I need your help with this integral

OpenStudy (adamaero):

x squared times... or ?

OpenStudy (bibby):

\[\int\limits_{}^{}\dfrac{x^2}{\sqrt{{1+x}}}dx\]

OpenStudy (dumbcow):

substitution u = 1+x du = dx

OpenStudy (dumbcow):

\[\int\limits_{}^{}\frac{(u-1)^2}{\sqrt{u}} du\]

OpenStudy (dumbcow):

\[= \int\limits \frac{u^2}{\sqrt{u}} -\frac{2u}{\sqrt{u}}+\frac{1}{\sqrt{u}}\]

OpenStudy (loser66):

another way: Let \(u =\sqrt{1+x}\), \(du=\dfrac{1}{2\sqrt{1+x}}dx\) and \(u^2= 1+x --> x=u^2-1\rightarrow x^2= (u^2-1)^2\) everything is ok from here.